y'' of y = x^-2 at x=2


1
Expert's answer
2021-11-01T08:46:29-0400

Let us find "y''" of "y = x^{-2}" at "x=2."


It follows that "y'=-2x^{-3},\ y''=(-3)(-2)x^{-4}=6x^{-4}."


Consequently, "y''(2)=6\cdot 2^{-4}=\frac{6}{2^4}=\frac{3}{2^3}=\frac{3}8."